OpenStudy (anonymous):

Help!! (3d²+2d-29) ÷ (d+3) Can you show me how to simlifly this by canceling like factors?

OpenStudy (anonymous):

Factor the expression on the left and factor the expression on the right. Then cancel like factors.

OpenStudy (anonymous):

Can the right be factored?

OpenStudy (amistre64):

the right cannot be factored

OpenStudy (amistre64):

IF the left side factors into something useful; it will have to be of the form: (d+3)(md + n)

OpenStudy (amistre64):

if we "foil" this out ... what would you say would be a smart value for m?

OpenStudy (anonymous):

@amistre64 Idk I tryed factoring it but 29 is prime with no like factors, and i still dont get how 2 foil.

OpenStudy (amistre64):

f1 l1 f2 l2 (d+3) (md + n) foil is a memory method for remembering how to keep track of what you are distributing thru. F = f1 f2 = md^2 O = f1 l2 = nd I = l1 f2 = 3md L = l1 l2 = 3n equating these to what we have md^2 = 3d^2 ; m=3 (n + 3m)d = 2d 3n = -29 ; n = -29/3 (-29/3 + 3(3))d = 2d ((-29+27)/3)d = 2d (-2/3)d = 2d is rather impossiible you sure you need to factor and cancel?

OpenStudy (anonymous):

All my teacher said 2 do was multiply or divide...

OpenStudy (amistre64):

if the remainder of the division is zero, then we know we can cancel out some like factors: a shortcut to seeing if the remainder is zero is to find a value for d that zeros out the bottom (d+3 = 0, when d=-3) and use it on the top to see if the top also zeros out.

OpenStudy (amistre64):

3d² + 2d - 29 3(9) + 2(-3) - 29 27 -6 - 29 is not zero, so the bottom is not a factor of the top

OpenStudy (anonymous):

ok, now what?

OpenStudy (amistre64):

if you are being asked to find the remainder ... 27-6-29 = -8 -8/(d+3) but it is not clear what your end game is spose to be

OpenStudy (amistre64):

prolly not

OpenStudy (amistre64):

no

OpenStudy (anonymous):

So.... what do we do then?

OpenStudy (amistre64):

there are over a hundred and sixty-three "things" we can do .... but without knowing what it is that you NEED to do (what is your goal for this), then it would be a bit futile to carry on

OpenStudy (amistre64):

IF your intent is to see what factors out, then we determined that at the start .... nothing factors out

OpenStudy (amistre64):

if all your teacher is asking you to do is divide .... then id suggest simply dividing out into a quotient and remainder

OpenStudy (amistre64):

3d - 7 --------------- d+3 | 3d^2+2d-29 -(3d^2+9d) --------- -7d-29 -(-7d-21) -------- -8